# chebgreen—: Learning and Interpolating Continuous Empirical Green’s Functions from Data

### v2

@@ Appendix A Proof for interpolating quasimatrices on tangent space of $H=(L^{2}(\Omega))^{K}$ space
$\operatorname{tr}\left(\Psi^{*}P_{\Phi}(\Gamma)\right)=\operatorname{tr}\left(\Psi^{*}\Gamma-\frac{1}{2}\Psi^{*}\left[\Phi(\Phi^{*}\Gamma+\Gamma^{*}\Phi)\right]\right).$
So it suffices to show that
$\operatorname{tr}\left(\left[\Phi(\Phi^{*}\Psi+\Psi^{*}\Phi)\right]^{*}\Gamma\right)=\operatorname{tr}\left(\Psi^{*}\left[\Phi(\Phi^{*}\Gamma+\Gamma^{*}\Phi)\right]\right).$ (20)
To see that left side of Eq. (A.1) satisfies
$\operatorname{tr}\left(\left[\Phi(\Phi^{*}\Psi+\Psi^{*}\Phi)\right]^{*}\Gamma\right)=\operatorname{tr}\left((\Phi^{*}\Psi+\Psi^{*}\Phi)(\Phi^{*}\Gamma)\right)$ (21)
consider the left side of Eq. (A.2),
$\operatorname{tr}\left([\Phi(\Phi^{*}\Psi+\Psi^{*}\Phi)]^{*}\Gamma\right)=\operatorname{tr}\left([\Phi(\Phi^{*}\Psi)]^{*}\Gamma+[\Phi(\Psi^{*}\Phi)]^{*}\Gamma\right),$
note that computing the term in the left of trace argument, $[\Phi(\Phi^{*}\Psi)]^{*}\Gamma,$ gives a $k\times k$ matrix $C$ such that
$C_{ij}=\left(\sum_{i=1}^{k}\Phi_{i}(\Phi_{j},\Psi_{k})_{L^{2}(\Omega)},\Gamma_{j}\right)_{L^{2}(\Omega)}.$
Expanding right side of Eq. (A.2),
$\operatorname{tr}\left((\Phi^{*}\Psi+\Psi^{*}\Phi)(\Phi^{*}\Gamma)\right)=\operatorname{tr}\left((\Phi^{*}\Psi)(\Phi^{*}\Gamma)+(\Psi^{*}\Phi)(\Phi^{*}\Gamma)\right)$
and multiplying out the term $(\Phi^{*}\Psi)(\Phi^{*}\Gamma)$, we find that this is equal to $C=[\Phi(\Phi^{*}\Psi)]^{*}\Gamma.$ Doing the same computation for the other term, we conclude that Eq. (A.2) holds.
Similarly, the right side of Eq. (A.1) gives
$\operatorname{tr}\left(\Psi^{*}\left[\Phi(\Phi^{*}\Gamma+\Gamma^{*}\Phi)\right]\right)=\operatorname{tr}\left((\Psi^{*}\Phi)(\Phi^{*}\Gamma+\Gamma^{*}\Phi)\right)$
Using the fact that $(\Xi^{*}\Omega)^{T}=\Omega^{*}\Xi$ for any $\Xi,\Omega\in H,$ which comes from the definition of the outer product on $H$ in Eq. (8), it follows that Eq. (A.1) holds. Hence $P_{\Phi}$ is an orthogonal projection.
Finally, to see that the image of $P_{\Phi}$ is all of $T_{\Phi}S_{K}(V),$ consider the operator $P^{\perp}_{\Phi}$ defined by the mapping $\Psi\mapsto\Phi\operatorname{sym}(\Phi^{*}\Psi).$ The normal space may be characterized as the set of all elements of $H$ of the form $\Phi M$ where $M$ is a symmetric matrix. Clearly any element of the normal space is the image of some $\Psi$ under this mapping. Since we have $\Psi=P_{\Phi}(\Psi)+P^{\perp}_{\Phi}(\Psi)$ for any $\Psi\in H,$ it follows that $P_{\Phi}$ is the orthogonal projection onto $T_{\Phi}S_{K}(H).$
Appendix B Test error for interpolation cases with noisy dataset

### v5

@@ 5 Conclusions
$\|r(t\eta)\|_{H}=O\bigl{(}t^{3}\|\eta\|_{H}^{3}\bigr{)}\quad\text{as }t\to 0.$
By (i) and (ii), we have
$R(\phi,0)=\phi,\quad DR(\phi,0)[\eta]=\eta.$
Thus,
$R(\phi,t\eta)=\phi+t\eta+\frac{t^{2}}{2}\,D^{2}R(\phi,0)[\eta,\eta]+r(t\eta).$
Subtracting $\phi+t\eta$ and taking the $H$-norm, we obtain
$\|R(\phi,t\eta)-(\phi+t\eta)\|_{H}\leq\frac{t^{2}}{2}\,\|D^{2}R(\phi,0)[\eta,\eta]\|_{H}+\|r(t\eta)\|_{H}.$
By assumption (iii),
$\|D^{2}R(\phi,0)[\eta,\eta]\|_{H}\leq L\,\|\eta\|_{H}^{2}.$
Hence,
$\|R(\phi,t\eta)-(\phi+t\eta)\|_{H}\leq\frac{L}{2}\,t^{2}\,\|\eta\|_{H}^{2}+\|r(t\eta)\|_{H}.$
Since $\|r(t\eta)\|_{H}=o(t^{2}\|\eta\|_{H}^{2})$, for all $\varepsilon_{0}>0$ there exists $\delta_{0}>0$ such that if $0<|t|<\delta_{0}$, then
$\|r(t\eta)\|_{H}\leq\varepsilon_{0}\,t^{2}\|\eta\|_{H}^{2}.$
Taking $\varepsilon_{0}=L/2$ yields
$\|r(t\eta)\|_{H}\leq\frac{L}{2}\,t^{2}\|\eta\|_{H}^{2}\quad\text{whenever }|t|<\delta_{0}.$
Thus, for $|t|<\delta$,
$\|R(\phi,t\eta)-(\phi+t\eta)\|_{H}\leq\frac{L}{2}\,t^{2}\,\|\eta\|_{H}^{2}+\frac{L}{2}\,t^{2}\,\|\eta\|_{H}^{2}=L\,t^{2}\,\|\eta\|_{H}^{2}.$
This completes the proof. ∎
Lemma 2.
Let
$H=\left(L^{2}(\Omega)\right)^{K}$
be a Hilbert space and let $S_{K}$ be the Hilbert Stiefel manifold of $H$. Define the map
$qf:U\subset H\to S_{K},$
by applying the Gram–Schmidt process, that is, define
$q_{1}(\phi_{1})=\frac{\phi_{1}}{\|\phi_{1}\|_{H}},$
and for $j=2,\dots,K$, recursively set
$p_{j}(\phi_{1},\dots,\phi_{j})=\phi_{j}-\sum_{i=1}^{j-1}\langle\phi_{j},q_{i}(\phi_{1},\dots,\phi_{i})\rangle_{H}\,q_{i}(\phi_{1},\dots,\phi_{i}),$
$q_{j}(\phi_{1},\dots,\phi_{j})=\frac{p_{j}(\phi_{1},\dots,\phi_{j})}{\|p_{j}(\phi_{1},\dots,\phi_{j})\|_{H}}.$
Then, the mapping
$qf(\phi)=\bigl{(}q_{1}(\phi_{1}),\;q_{2}(\phi_{1},\phi_{2}),\;\dots,\;q_{K}(\phi_{1},\dots,\phi_{K})\bigr{)}$
is smooth on the open set
$U=\{\phi\in H:\|\phi_{i}\|_{H}>0\text{ for }i=1,\dots,K,\text{ and }\{\phi_{1},\dots,\phi_{K}\}\text{ is linearly independent}\}.$
Proof.
We view the process as a composition of maps: First, the normalization $q_{1}(\phi_{1})=\phi_{1}/\|\phi_{1}\|_{H}$ is smooth on the set where $\|\phi_{1}\|_{H}>0$. For each $j\geq 2$, given that $q_{1},\dots,q_{j-1}$ are smooth, the map
$\phi_{j}\mapsto\operatorname{proj}_{\operatorname{span}\{\phi_{1},\dots,\phi_{j-1}\}}(\phi_{j})=\sum_{i=1}^{j-1}\langle\phi_{j},q_{i}\rangle_{H}\,q_{i},$
is smooth since the inner product and scalar multiplication are smooth operations in a Hilbert space. Then, subtracting this projection from $\phi_{j}$ is linear and bounded and hence smooth; the normalization step,
$q_{j}=\frac{p_{j}}{\|p_{j}\|_{H}},$
is smooth on the set of quasimatrices satisfying $p_{j}\neq 0$, which holds under the assumption that the input has full rank.
Since each $q_{j}$ is smooth and the overall mapping $qf$ is given by the composition of these maps, $qf$ is smooth on $U$. ∎
6 Test error for interpolation cases with noisy dataset